# Direct3D-S2: Gigascale 3D Generation Made Easy with Spatial Sparse Attention

### v1

@@ Abstract
Generating high-resolution 3D shapes using volumetric representations such as Signed Distance Functions (SDFs) presents substantial computational and memory challenges. We introduce Direct3D-S2, a scalable 3D generation framework based on sparse volumes that achieves superior output quality with dramatically reduced training costs. Our key innovation is the Spatial Sparse Attention (SSA) mechanism, which greatly enhances the efficiency of Diffusion Transformer (DiT) computations on sparse volumetric data. SSA allows the model to effectively process large token sets within sparse volumes, significantly reducing computational overhead and achieving a 3.9$\times$ speedup in the forward pass and a 9.6$\times$ speedup in the backward pass. Our framework also includes a variational autoencoder (VAE) that maintains a consistent sparse volumetric format across input, latent, and output stages. Compared to previous methods with heterogeneous representations in 3D VAE, this unified design significantly improves training efficiency and stability. Our model is trained on public available datasets, and experiments demonstrate that Direct3D-S2 not only surpasses state-of-the-art methods in generation quality and efficiency, but also enables training at 1024³ resolution using only 8 GPUs—a task typically requiring at least 32 GPUs for volumetric representations at $256^{3}$ resolution, thus making gigascale 3D generation both practical and accessible. Project page: https://nju-3dv.github.io/projects/Direct3D-S2/.
@@ 1 Introduction
Explicit latent methods have emerged as a compelling alternative to implicit ones, offering better interpretability, simpler training, and direct editing capabilities, while also adopting scalable architectures such as DiT [29]. For instance, Direct3D [38] proposes to use tri-plane latent representations to accelerate training and convergence. XCube [31] introduces hierarchical sparse voxel latent diffusion for $1024^{3}$ sparse volume generation, but only restricted to millions of valid voxels, limiting the final output quality. Trellis [39] integrates sparse voxel representations of $256^{3}$ resolution, with the rendering supervision for the VAE training. In general, due to high memory demands, existing explicit-latent methods are limited in output resolution. Scaling to $1024^{3}$ with sufficient latent tokens and valid voxels remains challenging, as the quadratic cost of full attention in DiT renders high-resolution training computationally prohibitive.
@@
To address the challenge of high-resolution 3D shape generation, we propose Direct3D-S2, a unified generative framework that utilizes sparse volumetric representations. At the core of our approach is a novel Spatial Sparse Attention (SSA) mechanism, which significantly improves the scalability of diffusion transformers in high-resolution 3D shape generation by selectively attending to spatially important tokens via learnable compression and selection modules. Specifically, we draw inspiration from the key principles of Native Sparse Attention (NSA) [45], which integrates compression, selection, and windowing to identify relevant tokens based on global-local interactions. While NSA is designed for structurally organized 1D sequences, it is not directly applicable to unstructured sparse 3D data. To adapt it, we redesign the block partitioning to preserve 3D spatial coherence and revise the core modules to accommodate the irregular nature of sparse volumetric tokens. This enables efficient processing of large token sets within sparse volumes. We implement a custom Triton [36] GPU kernel for SSA, achieving a 3.9$\times$ speedup in the forward pass and a 9.6$\times$ speedup in the backward pass compared to FlashAttention-2 at $1024^{3}$ resolution.
@@
Our framework also includes a VAE that maintains a consistent sparse volumetric format across input, latent, and output stages. This unified design eliminates the need for cross-modality translation—commonly seen in previous methods using mismatched representations such as point cloud input, 1D vector latent, and dense volume output—thereby significantly improving training efficiency, stability, and geometric fidelity. After the VAE training, the DiT with the proposed SSA will be trained on the converted latents, enabling scalable and efficient high-resolution 3D shape generation.
@@
Extensive experiments demonstrate that our approach successfully achieves high-quality and efficient gigascale 3D generation, a milestone previously unattainable by explicit 3D latent diffusion methods. Compared to prior native 3D diffusion techniques, our model consistently generates highly detailed 3D shapes while significantly reducing computational costs. Notably, Direct3D-S2 requires only 8 GPUs to train on public datasets [8, 9, 19] at a resolution of $1024^{3}$, in stark contrast to prior state-of-the-art methods, which typically require 32 or more GPUs even for training at $256^{3}$ resolution.
@@ 2.2 Large Scale 3D Latent Diffusion Model
Motivated by recent advances in Latent Diffusion Models (LDMs) [32] in 2D image generation, several methods have extended LDMs to 3D shape generation. These approaches broadly fall into two categories: vecset-based methods and voxel-based methods. Implicit vecset-based methods, such as 3DShape2Vecset [46], Michelangelo [49], CLAY [48], and CraftsMan3D [16], represent 3D shapes using latent vecset and reconstruct meshes through neural SDFs or occupancy fields. However, implicit methods are typically constrained by vecset size: larger vecset lead to more complex mappings to the 3D shape and require longer training times. In contrast, voxel-based methods, such as XCube [31] and Trellis [39], employs voxel grids as latent representations, providing more interpretability and easier training. Nevertheless, voxel-based methods face limitations in latent resolution due to cubic growth in GPU memory requirements and high computational costs associated with attention mechanisms. To address this issue, our work specifically targets reducing the computational overhead of attention mechanisms, thereby enabling the generation of high-resolution voxel-based latent representations that were previously infeasible.
@@ 3 Sparse SDF VAE
Figure 2: The framework of our Direct3D-S2. We propose a fully end-to-end sparse SDF VAE (SS-VAE), which employs a symmetric encoder-decoder network to efficiently encode high-resolution sparse SDF volumes into sparse latent representations $\mathbf{z}$. Then we train an image-conditioned diffusion transformer (SS-DiT) based on $\mathbf{z}$, and design a novel Spatial Sparse Attention (SSA) mechanism that significantly improves the training and inference efficiency of the DiT.
While variational autoencoders (VAEs) have become the cornerstone of 2D image generation by compressing pixel representations into compact latent spaces for efficient diffusion training, their extension to 3D geometry faces fundamental challenges. Unlike images with standardized pixel grids, 3D representations lack a unified structure – meshes, point clouds, and implicit fields each require specialized processing. This fragmentation forces existing 3D VAEs into asymmetric architectures with compromised efficiency. For instance, prominent approaches [6, 48, 50] based on vecset [46] encode the input point cloud into a vector set latent space before decoding it into SDF field, while Trellis [39] and XCube [31] rely on differentiable rendering or post-hoc neural kernel surface reconstruction [11] to bridge their latent spaces to usable meshes. These hybrid pipelines introduce computational bottlenecks and geometric approximations that limit their scalability to high-resolution 3D generation. In this paper, we propose a fully end-to-end sparse SDF VAE that employs a symmetric encoding-decoding network to encode high-resolution sparse SDF volumes into a sparse latent representation, significantly improving training efficiency while maintaining geometric precision.
@@ 3.1 Symmetric Network Architecture
Our fully end-to-end SDF VAE framework adopts a symmetric encoder-decoder network architecture, as illustrated in the upper half of Figure 2. Specifically, the encoder employs a hybrid framework combining sparse 3D convolution networks and transformer networks. We first extract local geometric features through a series of residual sparse 3D CNN blocks interleaved with 3D mean pooling operations, progressively downsampling the spatial resolution. We then process the sparse voxels as variable-length tokens and utilize transformer layers to capture global contextual information between the valid voxels. Inspired by Trellis [39], the feature of each valid voxel is augmented with positional encoding based on its 3D coordinates before being fed into multi-head self-attention layers. This hybrid design outputs sparse latent representations at reduced resolution $(\frac{R}{f})^{3}$, where $f$ denotes the downsampling factor. The decoder of our SS-VAE adopts a symmetric structure with respect to the encoder, leveraging transformer layers and sparse 3D CNN blocks to progressively upsample the latent representation and reconstruct the SDF volume $\tilde{V}$.
@@ 4.1 Spatial Sparse Attention
where $\mathbf{k}^{\text{cmp}}_{t}$ denotes the block-level key token, $\text{PE}(\cdot)$ is absolute position encoding, and $\delta(\cdot)$ represents operations of sparse 3D convolution and sparse 3D mean pooling. The sparse 3D compression module effectively captures block-level global information while significantly reducing the token count, thereby enhancing computational efficiency.
@@
where $\mathcal{B}_{\text{cmp}}$ denotes the set of compression blocks within the selection block, and $h_{s}$ represents the number of shared heads within a group. The top-$k$ selection blocks with the highest $\mathbf{s}^{\text{slc}}_{t}$ scores are selected, and all tokens contained within them are concatenated to form $\mathbf{k}^{\text{slc}}_{t}$ and $\mathbf{v}^{\text{slc}}_{t}$, which are used to compute the spatial blockwise selection attention. We implement the attention kernel using Triton [36], with a key challenge arising from the sparse 3D voxel structure where the number of tokens per block varies. To address this, we pre-compute the index $\mathcal{C}$ of the first token in each block as an input to the kernel and sort the input tokens based on their block indices, ensuring that all tokens belonging to the same block are contiguous in memory.
In our Triton-based implementation of the spatial blockwise selection attention kernel, two key challenges arise within sparse 3D voxel structures: 1) the number of tokens varies across different blocks, and 2) tokens within the same block may not be contiguous in HBM. To address these, we first sort the input tokens based on their block indices, then compute the starting index $\mathcal{C}$ of each block as kernel input. In the inner loop, $\mathcal{C}$ dynamically governs the loading of corresponding block tokens. The complete procedure of forward pass is formalized in Algorithm 1.
@@ 5.2 Implementation Details
DiT. Our SS-DiT comprises 24 layers of DiT blocks with a hidden dimension of 1024. We employ Grouped-Query Attention (GQA) [4] with a group number set to 2, where each group contains 16 attention heads. The hidden dimension of each head is configured as 32. For the spatial sparse attention (SSA) mechanism, we configure the resolution of the compression blocks to $m_{\text{cmp}}=4$ , the resolution of the selection blocks to $m_{\text{slc}}=8$, and the size of the sparse 3D windows $m_{\text{win}}=8$. We utilize DINO-v2 Large [28] to extract features from conditional images, with input images having a resolution of $518\times 518$. For the DiT, we implement a progressive training strategy that gradually increases the resolution from $256^{3}$ to $1024^{3}$ to accelerate convergence. Table 1 presents the average latent token count, learning rate, batch size, and training duration settings at different resolutions. We employ the AdamW optimizer and trained the model for a total of 7 days on 8 A100 GPUs. For the $1024^{3}$ resolution, we further filtered 68k high-fidelity 3D assets for training. Additionally, similar to Trellis [39], we trained an extra DiT to predict the indices of the sparse latent tokens $\mathbf{z}$, which took 7 days on 8 A100 GPUs.
Table 1: Training configurations for our DiT at four voxel resolutions. Res., TC, LR, BS, and TT denote resolution, token count, learning rate, batch size, and total training time, respectively.
Res. TC LR BS TT
@@ 6.1 Ablation Studies
Image-to-3D Generation in Different Resolution. We present the generation results of our Direct3D-S2 across four resolutions {$256^{3},384^{3},512^{3},1024^{3}$} in Figure 8. The results demonstrate that increasing resolution progressively improves mesh quality. At lower resolutions $256^{3}$ and $384^{3}$, the generated meshes exhibit limited geometric details and misalignment with input images. At $512^{3}$ resolution, the meshes display significantly enhanced high-frequency geometric details. Further increasing the resolution to $1024^{3}$ yields meshes with sharper edges and improved alignment with input image details.
@@
Runtime of Different Attention Mechanisms. We implemented a custom Triton [36] GPU kernel for SSA. And we compare the forward and backward execution times of our SSA with those of FlashAttention-2 [7] across various token counts, using the implementation from Xformers [14] for FlashAttention-2. The comparison results are shown in Figure 7, which indicate that our SSA achieves comparable speeds to FlashAttention-2 when the token count is low; however, as the token count increases, the speed advantage of our SSA becomes more pronounced. Specifically, when the token count reaches 128k, the forward and backward speeds of our SSA are 3.9×3.9$\times$3.9 × and 9.6×9.6$\times$9.6 × faster than those of FlashAttention-2, respectively, demonstrating the efficiency of our proposed SSA.
Effectiveness of SSA. We conduct ablation studies to validate the robustness of SSA. Given the insufficient geometric fidelity at $256^{3}$/ $384^{3}$ resolutions, which do not adequately reflect the model’s precision, and prohibitive computational costs at $1024^{3}$ resolution, we perform experiments at $512^{3}$ resolution. We establish three comparative configurations: 1) Full attention: directly training the DiT with full attention proves to be inefficient. Therefore, following Trellis’ latent packing strategy [39], we group latent tokens within $2^{3}$ local regions to reduce the token count before feeding them into the DiT blocks. 2) NSA: process latent tokens as 1D sequences with fixed-length block partitioning, disregarding spatial coherence. 3) Our proposed SSA. The qualitative results are illustrated in Figure 10. It is evident that the full attention variant produces meshes with high-frequency surface artifacts, attributed to its forced packing operation that disrupts local geometric continuity. The NSA implementation exhibits training instability due to positional ambiguity in block partitioning, resulting in less smooth meshes. In contrast, our SSA not only preserves the details of the meshes, but also yields a smoother and more organized surface, thereby demonstrating the effectiveness of our proposed SSA mechanism.
Effect of Sparse Conditioning Mechanism. We perform ablation experiments to validate the effect of the sparse conditioning mechanism at $512^{3}$ resolution. As demonstrated in Figure 11, the exclusion of non-foreground conditioning tokens through sparse conditioning enables the generated meshes to achieve significantly better alignment with the input images.

### v2

@@ Abstract
Generating high-resolution 3D shapes using volumetric representations such as Signed Distance Functions (SDFs) presents substantial computational and memory challenges. We introduce Direct3D-S2, a scalable 3D generation framework based on sparse volumes that achieves superior output quality with dramatically reduced training costs. Our key innovation is the Spatial Sparse Attention (SSA) mechanism, which greatly enhances the efficiency of Diffusion Transformer (DiT) computations on sparse volumetric data. SSA allows the model to effectively process large token sets within sparse volumes, substantially reducing computational overhead and achieving a 3.9$\times$ speedup in the forward pass and a 9.6$\times$ speedup in the backward pass. Our framework also includes a variational autoencoder (VAE) that maintains a consistent sparse volumetric format across input, latent, and output stages. Compared to previous methods with heterogeneous representations in 3D VAE, this unified design significantly improves training efficiency and stability. Our model is trained on public available datasets, and experiments demonstrate that Direct3D-S2 not only surpasses state-of-the-art methods in generation quality and efficiency, but also enables training at 1024³ resolution using only 8 GPUs, a task typically requiring at least 32 GPUs for volumetric representations at $256^{3}$ resolution, thus making gigascale 3D generation both practical and accessible. Project page: https://www.neural4d.com/research/direct3d-s2.
@@ 1 Introduction
Explicit latent methods have emerged as a compelling alternative to implicit ones, offering better interpretability, simpler training, and direct editing capabilities, while also adopting scalable architectures such as DiT [30]. For instance, Direct3D [39] proposes to use tri-plane latent representations to accelerate training and convergence. XCube [32] introduces hierarchical sparse voxel latent diffusion for $1024^{3}$ sparse volume generation, but only restricted to millions of valid voxels, limiting the final output quality. Trellis [40] integrates sparse voxel representations of $256^{3}$ resolution, with the rendering supervision for the VAE training. In general, due to high memory demands, existing explicit latent methods are limited in output resolution. Scaling to $1024^{3}$ with sufficient latent tokens and valid voxels remains challenging, as the quadratic cost of full attention in DiT renders high-resolution training computationally prohibitive.
@@
To address the challenge of high-resolution 3D shape generation, we propose Direct3D-S2, a unified generative framework that utilizes sparse volumetric representations. At the core of our approach is a novel Spatial Sparse Attention (SSA) mechanism, which substantially improves the scalability of diffusion transformers in high-resolution 3D shape generation by selectively attending to spatially important tokens via learnable compression and selection modules. Specifically, we draw inspiration from the key principles of Native Sparse Attention (NSA) [46], which integrates compression, selection, and windowing to identify relevant tokens based on global-local interactions. While NSA is designed for structurally organized 1D sequences, it is not directly applicable to unstructured, sparse 3D data. To adapt it, we redesign the block partitioning to preserve 3D spatial coherence and revise the core modules to accommodate the irregular nature of sparse volumetric tokens. This enables efficient processing of large token sets within sparse volumes. We implement a custom Triton [37] GPU kernel for SSA, achieving a 3.9$\times$ speedup in the forward pass and a 9.6$\times$ speedup in the backward pass compared to FlashAttention-2 at $1024^{3}$ resolution.
@@
Our framework also includes a VAE that maintains a consistent sparse volumetric format across input, latent, and output stages. This unified design eliminates the need for cross-modality translation, commonly seen in previous methods using mismatched representations such as point cloud input, 1D vector latent, and dense volume output, thereby improving training efficiency, stability, and geometric fidelity. After the VAE training, the DiT with the proposed SSA will be trained on the converted latents, enabling scalable and efficient high-resolution 3D shape generation.
@@
Extensive experiments demonstrate that our approach successfully achieves high-quality and efficient gigascale 3D generation, a milestone previously unattainable by explicit 3D latent diffusion methods. Compared to prior native 3D diffusion techniques, our model consistently generates highly detailed 3D shapes while considerably reducing computational costs. Notably, Direct3D-S2 requires only 8 GPUs to train on public datasets [8, 9, 20] at a resolution of $1024^{3}$, in stark contrast to prior state-of-the-art methods, which typically require 32 or more GPUs even for training at $256^{3}$ resolution.
@@ 2.2 Large Scale 3D Latent Diffusion Model
Motivated by recent advances in Latent Diffusion Models (LDMs) [33] in 2D image generation, several methods have extended LDMs to 3D shape generation. These approaches broadly fall into two categories: vecset-based methods and voxel-based methods. Implicit vecset-based methods, such as 3DShape2Vecset [47], Michelangelo [50], CLAY [49], and CraftsMan3D [17], represent 3D shapes using latent vecset and reconstruct meshes through neural SDFs or occupancy fields. However, implicit methods are typically constrained by the size of vecset: larger vecset leads to more complex mappings to the 3D shape and requires longer training times. In contrast, voxel-based methods, such as XCube [32], Trellis [40], and more recent works [11, 45], employ voxel grids as latent representations, providing more interpretability and easier training. Nevertheless, voxel-based methods face limitations in latent resolution due to cubic growth in GPU memory requirements and high computational costs associated with attention mechanisms. To address this issue, our work specifically targets reducing the computational overhead of attention mechanisms, thereby enabling the generation of high-resolution voxel-based latent representations that were previously infeasible.
@@ 3 Sparse SDF VAE
Figure 2: The framework of our Direct3D-S2. We propose a fully end-to-end sparse SDF VAE (SS-VAE), which employs a symmetric encoder-decoder network to efficiently encode high-resolution sparse SDF volumes into sparse latent representations $\mathbf{z}$. Then we train an image-conditioned diffusion transformer (SS-DiT) based on $\mathbf{z}$, and design a novel Spatial Sparse Attention (SSA) mechanism that substantially improves the training and inference efficiency of the DiT.
While variational autoencoders (VAEs) have become the cornerstone of 2D image generation by compressing pixel representations into compact latent spaces for efficient diffusion training, their extension to 3D geometry faces fundamental challenges. Unlike images with standardized pixel grids, 3D representations lack a unified structure, such as meshes, point clouds, and implicit fields, each requires specialized processing. This fragmentation forces existing 3D VAEs into asymmetric architectures with compromised efficiency. For instance, prominent approaches [6, 49, 51] based on vecset [47] encode the input point cloud into a vector set latent space before decoding it into SDF field, while Trellis [40] and XCube [32] rely on differentiable rendering or neural kernel surface reconstruction [12] to bridge their latent spaces to usable meshes. These hybrid pipelines introduce computational bottlenecks and geometric approximations that limit their scalability to high-resolution 3D generation. In this paper, we propose a fully end-to-end sparse SDF VAE that employs a symmetric encoding-decoding network to encode high-resolution sparse SDF volumes into a sparse latent representation, substantially improving training efficiency while maintaining geometric precision.
@@ 3.1 Symmetric Network Architecture
Our fully end-to-end SDF VAE framework adopts a symmetric encoder-decoder network architecture, as illustrated in the upper half of Figure 2. Specifically, the encoder employs a hybrid framework combining sparse 3D convolution networks and transformer networks. We first extract local geometric features through a series of residual sparse 3D CNN blocks interleaved with 3D mean pooling operations, progressively downsampling the spatial resolution. We then process the sparse voxels as variable-length tokens and utilize shifted window attention to capture local contextual information between the valid voxels. Inspired by Trellis [40], the feature of each valid voxel is augmented with positional encoding based on its 3D coordinates before being fed into 3D shift window attention layers. This hybrid design outputs sparse latent representations at reduced resolution $(\frac{R}{f})^{3}$, where $f$ denotes the downsampling factor. The decoder of our SS-VAE adopts a symmetric structure with respect to the encoder, leveraging attention layers and sparse 3D CNN blocks to progressively upsample the latent representation and reconstruct the SDF volume $\tilde{V}$.
@@ 4.1 Spatial Sparse Attention
where $\mathbf{k}^{\text{cmp}}_{t}$ denotes the block-level key token, $\text{PE}(\cdot)$ is absolute position encoding, and $\delta(\cdot)$ represents operations of sparse 3D convolution and sparse 3D mean pooling. The sparse 3D compression module effectively captures block-level global information while reducing the number of tokens, thereby enhancing computational efficiency.
@@
where $\mathcal{B}_{\text{cmp}}$ denotes the set of compression blocks within the selection block, and $h_{s}$ represents the number of shared heads within a group. The top-$k$ selection blocks with the highest $\mathbf{s}^{\text{slc}}_{t}$ scores are selected, and all tokens contained within them are concatenated to form $\mathbf{k}^{\text{slc}}_{t}$ and $\mathbf{v}^{\text{slc}}_{t}$, which are used to compute the spatial blockwise selection attention.
We implement the spatial blockwise selection attention kernel using Triton [37], with two key challenges arising from the sparse 3D voxel structures: 1) the number of tokens varies across different blocks, and 2) tokens within the same block may not be contiguous in HBM. To address these, we first sort the input tokens based on their block indices, then compute the starting index $\mathcal{C}$ of each block as kernel input. In the inner loop, $\mathcal{C}$ dynamically governs the loading of corresponding block tokens. The complete procedure of forward pass is formalized in Algorithm 1.
@@ 5.2 Implementation Details
DiT. Our SS-DiT comprises 24 layers of DiT blocks with a hidden dimension of 1024. We employ Grouped-Query Attention (GQA) [4] with a group number set to 2, where each group contains 16 attention heads. The hidden dimension of each head is configured as 32. For the spatial sparse attention (SSA) mechanism, we configure the resolution of the compression blocks to $m_{\text{cmp}}=4$ , the resolution of the selection blocks to $m_{\text{slc}}=8$, and the size of the sparse 3D windows $m_{\text{win}}=8$. We utilize DINO-v2 Large [29] to extract features from conditional images, with input images having a resolution of $518\times 518$. For the DiT, we implement a progressive training strategy that gradually increases the resolution from $256^{3}$ to $1024^{3}$ to accelerate convergence. Table 1 presents the average number of latent tokens, learning rate, batch size, and training duration settings at different resolutions. We employ the AdamW optimizer and trained the model for a total of 7 days on 8 A100 GPUs. For the $1024^{3}$ resolution, we further filtered 68k high-fidelity 3D assets for training. Additionally, similar to Trellis [40], we trained an extra DiT to predict the indices of the sparse latent tokens $\mathbf{z}$, which took 7 days on 8 A100 GPUs.
Table 1: Training configurations for DiT at four voxel resolutions. Res., NT, LR, BS, and TT denote resolution, number of tokens, learning rate, batch size, and total training time, respectively.
Res. NT LR BS TT
@@ 6.1 Ablation Studies
Image-to-3D Generation in Different Resolution. We present the generation results of our Direct3D-S2 across four resolutions {$256^{3},384^{3},512^{3},1024^{3}$} in Figure 8. The results demonstrate that increasing resolution progressively improves mesh quality. At lower resolutions $256^{3}$ and $384^{3}$, the generated meshes exhibit limited geometric details and misalignment with input images. At $512^{3}$ resolution, the meshes display enhanced high-frequency geometric details. Further increasing the resolution to $1024^{3}$ yields meshes with sharper edges and improved alignment with input image details.
@@
Runtime of Different Attention Mechanisms. We implemented a custom Triton [37] GPU kernel for SSA. And we compare the forward and backward execution times of our SSA with those of FlashAttention-2 [7] across various number of tokens, using the implementation from Xformers [15] for FlashAttention-2. The comparison results are shown in Figure 7, which indicate that our SSA achieves comparable speeds to FlashAttention-2 when the number of tokens is low; however, as the number of tokens increases, the speed advantage of our SSA becomes more pronounced. Specifically, when the number of tokens reaches 128k, the forward and backward speeds of our SSA are 3.9×3.9$\times$3.9 × and 9.6×9.6$\times$9.6 × faster than those of FlashAttention-2, respectively, demonstrating the efficiency of our proposed SSA.
Effectiveness of SSA. We conduct ablation studies to validate the robustness of SSA. Given the insufficient geometric fidelity at $256^{3}$/ $384^{3}$ resolutions, which do not adequately reflect the model’s precision, and prohibitive computational costs at $1024^{3}$ resolution, we perform experiments at $512^{3}$ resolution. We establish three comparative configurations: 1) Full attention: directly training the DiT with full attention proves to be inefficient. Therefore, following Trellis’ latent packing strategy [40], we group latent tokens within $2^{3}$ local regions to reduce the number of tokens before feeding them into the DiT blocks. 2) NSA: process latent tokens as 1D sequences with fixed-length block partitioning, disregarding spatial coherence. 3) Our proposed SSA. The qualitative results are illustrated in Figure 10. It is evident that the full attention variant produces meshes with high-frequency surface artifacts, attributed to its forced packing operation that disrupts local geometric continuity. The NSA implementation exhibits training instability due to positional ambiguity in block partitioning, resulting in less smooth meshes. In contrast, our SSA not only preserves the details of the meshes, but also yields a smoother and more organized surface, thereby demonstrating the effectiveness of our proposed SSA mechanism.
Effect of Sparse Conditioning Mechanism. We perform ablation experiments to validate the effect of the sparse conditioning mechanism at $512^{3}$ resolution. As demonstrated in Figure 11, the exclusion of non-foreground conditioning tokens through sparse conditioning enables the generated meshes to achieve notably better alignment with the input images.